# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 6.5 Properties of the Dual Polytope
    The inner radius of $Q$ is at least $\frac{1}{45}$ when centered at $\mathbf{0}$;
@@
    The outer radius of $Q$ is at most $1$ when centered at $\mathbf{0}$.
Lemma 56.
For any $k\in\mathbb{N}$, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{k+5}$ and $n=4k+7$ points $a_{1},\ldots,a_{n}\in\mathbb{B}_{1}^{k+5}$ such that $Q:=\operatorname{conv}(a_{1},\ldots,a_{n})$ satisfies
@@
denote the polytope obtained from shifting its center $(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$ to $\mathbf{0}_{d}$. Here, as in Lemma 55,
@@
By applying row rescalings we transform the constraints (22 - 27) into a matrix $A\in\mathbb{R}^{(4k+7)\times(k+5)}$ such that
@@
Then by Lemma 54 and 9, the inner and outer ball of $\tilde{Q}$ satisfy
@@
Also, by Lemma 51, 10 and 9, the inner and outer ball of $\tilde{Q}\cap W$ satisfy
@@
The lemma then follows from taking $Q=\frac{1}{30}\tilde{Q}$. ∎
@@ 6.6 Perturbation Analysis and Proof of the Lower Bound
In this subsection, we study the number of edges of the intersection polygon $Q\cap W$ after perturbation and prove our main lower bound theorem (Theorem 49). To show that our construction has many edges even after perturbation, we require the following two statements:
Lemma 57.
@@
Now the desired result follows by $\bar{Q}\subseteq Q+\varepsilon\mathbb{B}_{1}^{d}\subseteq Q+\frac{\varepsilon}{r/2}Q$ and the fact that $(1+x)^{-1}\geq 1-x$ for every $x>-1$. ∎
Lemma 58.
@@
Without loss of generality, re-scale $T$ so that $\mathbb{B}_{2}^{2}\subseteq T\subseteq(1+\varepsilon)\cdot\mathbb{B}_{2}^{2}$, where $\varepsilon=\beta/\alpha-1>0$.
Consider any edge $[q_{1},q_{2}]\subseteq T$ and let $p\in[q_{1},q_{2}]$ denote the minimum-norm point in this edge. Then we have $\|q_{1}-p\|^{2}=\|q_{1}\|^{2}+\|p\|^{2}-2\langle q_{1},p\rangle$. Since $p$ is the minimum-norm point, we have $\langle q_{1},p\rangle\geq\|p\|^{2}$, and hence $\|q_{1}-p\|^{2}\leq\|q_{1}\|^{2}-\|p\|^{2}\leq(1+\varepsilon)^{2}-\|p\|^{2}$. Since $p$ lies on the boundary of $T$ we have $\|p\|\geq 1$, which implies that $\|q_{1}-p\|^{2}\leq(1+\varepsilon)^{2}-1=2\varepsilon+\varepsilon^{2}$. The analogous argument for $\|q_{2}-p\|$ and the triangle inequality tell us that $\|q_{1}-q_{2}\|\leq 2\sqrt{2\varepsilon+\varepsilon^{2}}\leq 4\sqrt{\varepsilon}$. The choice of the edge $[q_{1},q_{2}]$ was arbitrary, hence every edge of $T$ has length as most $4\sqrt{\varepsilon}$.
But $T$ has perimeter at least $2\pi$. Since the perimeter of a polygon is equal to the sum of the lengths of its edges, this implies that $T$ has at least $\frac{2\pi}{4\sqrt{\varepsilon}}>\frac{1}{\sqrt{\varepsilon}}$ edges. ∎
Now, we can prove our generic lower bound Theorem 50 on the shadow size under adversarial $\ell_{1}$-perturbations.
Proof of Theorem 50.
Fix any $d>5$, let $k=d-5$ and observe that $n=4k+7$. Let $\bar{a}_{1},\ldots,\bar{a}_{n}$ be as constructed in Lemma 56. Then we have
@@
For any set of points $a_{1},\ldots,a_{n}$ such that $\|a_{i}-\bar{a}_{i}\|_{1}\leq\varepsilon$ for each $i\in[n]$, by Lemma 57 we have
@@
Therefore, we can bound the inner and outer radius of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ by
$\frac{\cos(\pi/2^{k+2})}{30\cdot(1+90\varepsilon)}\cdot\mathbb{B}_{2}^{k+5}\cap W\subseteq\operatorname{conv}(a_{1},\ldots,a_{n})\cap W\subseteq\cdot\frac{1}{30\cdot(1-90\varepsilon)}\cdot\mathbb{B}_{2}^{k+5}\cap W.$
It follows from Lemma 58 that the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ has at least $\left(\frac{1+90\varepsilon}{(1-90\varepsilon)\cos(\pi/2^{k+2})}-1\right)^{1/2}\geq$ $\Omega(\frac{1}{\sqrt{\varepsilon+4^{-k}}})$ edges. ∎
@@
Proof of Theorem 49.
Using concentration of Gaussian distribution in 13, we find that if $\sigma\leq 1/(360d\sqrt{\log n})$, then with probability at least $1-\binom{n}{d}^{-1}$, we have $\max_{i\in[n]}\|a_{i}-\bar{a}_{i}\|_{2}\leq 4\sigma\sqrt{d\log n}\leq\frac{1}{90\sqrt{d}}$. The result follows from Theorem 50 and the fact that $\|x\|_{1}\leq\sqrt{d}\|x\|_{2}$ for every $x\in\mathbb{R}^{d}$. ∎
@@ 6.7 Experimental Results
To measure whether analysis in Theorem 49 is tight or not, we ran numerical experiments. Using Python and Gurobi 10.0.3, we constructed a matrix $A$ such that
@@
as described earlier in this section. Writing $R$ as the maximum Euclidean norm among the row vectors of $A$, we sampled $\hat{A}$ with independent Gaussian distributed entries with standard deviation $\sigma R$ and $\mathbb{E}[\hat{A}]=A$. To approximate the shadow size, we optimized the objective vectors $\cos(\frac{(i+0.3)\pi}{2^{k+4}})x+\sin(\frac{(i+0.3)\pi}{2^{k+4}})y$, with $i=0,\dots,2^{k+5}-1$, over the polyhedron $\{z\in\mathbb{R}^{k+5}:\hat{A}z\leq 1\}$ and counted the number of distinct values $(x,y)$ found among the solutions. Two consecutive solutions were counted as distinct if their $x,y$ coordinates differed in $\ell_{1}$ norm by at least $10^{-11}$. [^2] The code used for this experiment is available at https://github.com/sophiehuiberts/shadow-size/tree/main/HLZ23
When $\sigma=0$, our code found $2^{k+1}$ such distinct points. For $\sigma>0$, Theorem 49 shows that we expect to find at least $\Omega\Big{(}\min\Big{(}\frac{1}{\sqrt{d\sigma\sqrt{\log d}}},2^{k}\Big{)}\Big{)}$ distinct pairs $(x,y)$.

### v3

@@ 6.4 Properties of the Polar Polytope
    The inner $\ell_{1}$-radius of $Q$ is at least $\frac{1}{45}$ when centered at $\mathbf{0}$;
@@
    The outer $\ell_{1}$-radius of $Q$ is at most $1$ when centered at $\mathbf{0}$.
Lemma 6.12.
For any $k\in\mathbb{N}$, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{k+5}$ and $n=4k+7$ points $\bar{a}_{1},\ldots,\bar{a}_{n}\in\mathbb{B}_{1}^{k+5}$ such that $Q:=\operatorname{conv}(\bar{a}_{1},\ldots,\bar{a}_{n})$ satisfies
@@
denote the polytope obtained from shifting its center $(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$ to $\mathbf{0}_{d}$. Here, as in Lemma 6.9,
@@
By applying row rescalings we transform the constraints (24 - 29) into a matrix $A\in\mathbb{R}^{(4k+7)\times(k+5)}$ such that
@@
Then by Lemma 6.8 and Fact 2, the inner and outer ball of $\tilde{Q}$ satisfy
@@
Also, by Lemma 6.3, Fact 3 and Fact 2, the inner and outer ball of $\tilde{Q}\cap W$ satisfy
@@
The lemma then follows from taking $Q=\frac{1}{30}\tilde{Q}$.
@@ 6.5 Perturbation Analysis and Proof of the Lower Bound
In this subsection, we study the number of edges of the intersection polygon $Q\cap W$ after perturbation and prove our main lower bound theorem (Theorem 6.1). To show that our construction has many edges even after perturbation, we require the following two statements:
Lemma 6.14.
@@
Note that $1-x^{2}\leq 1$, so $1-x\leq 1/(1+x)$ for any $1+x>0$. Combining with $\bar{Q}\subseteq Q+\varepsilon\mathbb{B}_{1}^{d}\subseteq Q+\frac{\varepsilon}{r/2}Q$ we conclude the desired result.
Lemma 6.16.
@@
Without loss of generality, re-scale $T$ so that $\mathbb{B}_{2}^{2}\subseteq T\subseteq(1+\varepsilon)\cdot\mathbb{B}_{2}^{2}$, where $\varepsilon=\beta/\alpha-1>0$. Note that since $\beta\leq 2\alpha$ we have $\varepsilon\leq 1$.
Consider any edge $q_{1},q_{2}]\subseteq T$ and let $p\in[q_{1},q_{2}]$ denote the minimum-norm point in this edge. Then we have $\|q_{1}-p\|^{2}=\|q_{1}\|^{2}+\|p\|^{2}-2\langle q_{1},p\rangle$. Since $p$ is the minimum-norm point, we have $\langle q_{1},p\rangle\geq\|p\|^{2}$, and hence $\|q_{1}-p\|^{2}\leq\|q_{1}\|^{2}-\|p\|^{2}\leq(1+\varepsilon)^{2}-\|p\|^{2}$. Since $p$ lies on the boundary of $T$ we have $\|p\|\geq 1$, which implies that $\|q_{1}-p\|^{2}\leq(1+\varepsilon)^{2}-1=2\varepsilon+\varepsilon^{2}$. The analogous argument for $\|q_{2}-p\|$ and the triangle inequality tell us that $\|q_{1}-q_{2}\|\leq 2\sqrt{2\varepsilon+\varepsilon^{2}}\leq 4\sqrt{\varepsilon}$, where we use $\varepsilon\leq 1$ at the second step. The choice of the edge $q_{1},q_{2}]$ was arbitrary, hence every edge of $T$ has length as most $4\sqrt{\varepsilon}$.
But $T$ has perimeter at least $2\pi$. Since the perimeter of a polygon is equal to the sum of the lengths of its edges, this implies that $T$ has at least $\frac{2\pi}{4\sqrt{\varepsilon}}>\frac{1}{\sqrt{\varepsilon}}$ edges.
Now, we can prove the generic lower bound claimed in Theorem 6.2 on the shadow size under adversarial $\ell_{1}$-perturbations.
Proof 6.18 (Proof of Theorem 6.2).
Fix any $d>5$, let k=d5k=d-5 and observe that $n=4k+7$. Let $\bar{a}_{1},\ldots,\bar{a}_{n}$ be as constructed in Lemma 6.12. Then we have
@@
For any set of points $a_{1},\ldots,a_{n}$ such that $\|a_{i}-\bar{a}_{i}\|_{1}\leq\varepsilon$ for each $i\in[n]$. By Lemma 6.14 and setting $r=1/45$, we have
$\operatorname{conv}(\bar{a}_{1},\ldots,\bar{a}_{n})\subseteq\frac{1}{1-2\varepsilon/r}\operatorname{conv}(a_{1},\ldots,a_{n})=\frac{1}{1-90\varepsilon}\operatorname{conv}(a_{1},\ldots,a_{n}).$
Therefore,
@@
Therefore, we can bound the inner and outer $\ell_{1}$-radius of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ by
$\frac{(1-90\varepsilon)\cdot\cos(\pi/2^{k+2})}{30}\cdot\mathbb{B}_{2}^{k+5}\cap W\subseteq\operatorname{conv}(a_{1},\ldots,a_{n})\cap W\subseteq\cdot\frac{1+45\varepsilon}{30}\cdot\mathbb{B}_{2}^{k+5}\cap W.$
It then follows from Lemma 6.16 that the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ has at least
$\left(\frac{(1-90\varepsilon)\cdot\cos(\pi/2^{k+2})}{(1+45\varepsilon)-(1-90\varepsilon)\cdot\cos(\pi/2^{k+2})}\right)^{1/2}$
$\geq$~ $\left(\frac{(1-90\varepsilon)\cdot(1-\frac{\pi^{2}}{4^{k+2}})}{(1+45\varepsilon)-(1-90\varepsilon)\cdot(1-\frac{\pi^{2}}{4^{k+2}})}\right)^{1/2}$ (By $\cos(x)\geq 1-x^{2}$ for $x\leq\frac{\pi}{2}$)
$\geq$~ $\left(\frac{1-90\varepsilon-\frac{\pi^{2}}{4^{k+2}}}{(1+45\varepsilon)-(1-90\varepsilon-\frac{\pi^{2}}{4^{k+2}})}\right)^{1/2}$
$\geq$~ $\Omega\left(\frac{1}{\sqrt{\varepsilon+4^{-k}}}\right)$
edges.
@@
Proof 6.19 (Proof of Theorem 6.1).
Using concentration of Gaussian distribution in Corollary 2.6, we find that if $\sigma\leq 1/(360d\sqrt{\log n})$, then with probability at least $1-\binom{n}{d}^{-1}$, we have $\max_{i\in[n]}\|a_{i}-\bar{a}_{i}\|_{2}\leq 4\sigma\sqrt{d\log n}\leq\frac{1}{90\sqrt{d}}$. The result follows from Theorem 6.2 and the fact that $\|x\|_{1}\leq\sqrt{d}\|x\|_{2}$ for every $x\in\mathbb{R}^{d}$.
@@ 6.6 Experimental Results
To measure whether the analysis in Theorem 6.1 is tight or not, we ran numerical experiments. Using Python and Gurobi 10.0.3, we constructed a matrix $A$ such that
@@
as described earlier in this section. Writing $R$ as the maximum Euclidean norm among the row vectors of $A$, we sampled $\hat{A}$ with independent Gaussian distributed entries with standard deviation $\sigma R$ and $\mathbb{E}[\hat{A}]=A$. To approximate the shadow size, we optimized the objective vectors $\cos(\frac{(i+0.3)\pi}{2^{k+4}})x+\sin(\frac{(i+0.3)\pi}{2^{k+4}})y$, with $i=0,\dots,2^{k+5}-1$, over the polyhedron $\{z\in\mathbb{R}^{k+5}:\hat{A}z\leq 1\}$ and counted the number of distinct values $(x,y)$ found among the solutions. Two consecutive solutions were counted as distinct if their $x,y$ coordinates differed in $\ell_{1}$ norm by at least $10^{-11}$. [^3] The code used for this experiment is available at https://github.com/sophiehuiberts/shadow-size/tree/main/HLZ23 Note that the number of vertices of this projection of the perturbed auxiliary polytope is equal to the number of vertices of the section of the perturbed polar polytope, which then describes the shadow size.
When $\sigma=0$, our code found $2^{k+1}$ such distinct points. For $\sigma>0$, Theorem 6.1 shows that we expect to find at least $\Omega\Big(\min\Big(\frac{1}{\sqrt{d\sigma\sqrt{\log d}}},2^{k}\Big)\Big)$ distinct pairs $(x,y)$.